OpenStudy (anonymous):

A box of groceries requires 5.0 newtons of force to lift it up 1.0 meter. How much work is done? 0.0 J 1.0 J 2.0 J 5.0 J

hartnn (hartnn):

we just did this ?

hartnn (hartnn):

W = F*d = 5*1

OpenStudy (anonymous):

5.0 J

OpenStudy (anonymous):

i ment this one sorry How much work is done when a force of 15.0 newtons is used to lift a box 3.0 meters into the air? 0.0 J 3.0 J 15.0 J 45 J

hartnn (hartnn):

same formula W =F*d = 15*3 =... ?

OpenStudy (anonymous):

45:)

hartnn (hartnn):

yes

OpenStudy (anonymous):

another medal for you :)